It is possible for a female aged 31 to buy a certain insurance policy by paying $600 a year for 10 years certain or $X a year for 5 years (but only if the policyholder is alive). Find X, assuming j1=8%.
1
Expert's answer
2019-05-16T10:18:22-0400

If we use the formula for compound interest, then:

600×(1 + 0.08)^10 = X×(1 + 0.08)^5,

X = 600×(1 + 0.08)^5 = 881.6.